OpenStudy (anonymous):

What is the value of x? x + 12 = -7

OpenStudy (alexandervonhumboldt2):

subtract 12 from both sides

OpenStudy (alexandervonhumboldt2):

x+12-12=-7-12 x=-7-12=?

OpenStudy (anonymous):

4x=12-8 4x=4 x=4/4 x=1 answer is (1,12

OpenStudy (alexandervonhumboldt2):

what? @terron_jones2003

OpenStudy (anonymous):

I think its -19

OpenStudy (anonymous):

that is for your first question

OpenStudy (anonymous):

right??

OpenStudy (alexandervonhumboldt2):

@terron_jones2003 your solving is incorrect

OpenStudy (alexandervonhumboldt2):

@paula. you are correct

OpenStudy (anonymous):

im wrong

OpenStudy (anonymous):

Nice, thanks! @AlexandervonHumboldt2